OpenStudy (lena772):

If 78.4 grams of aluminum metal react with excess hydrochloric acid, how many grams of hydrogen gas can be produced? 2 Al + 6 HCl “yields”/ 2 AlCl3 + 3 H2 17.6 g H2 8.79 g H2 3.91 g H2 2.93 g H2

OpenStudy (anonymous):

Well do you know what the molar mas of aluminum is?

OpenStudy (anonymous):

mass*

OpenStudy (lena772):

26.98 g

OpenStudy (anonymous):

Yes. So, in the equation. The numbers before the compounds are mole ratios. For every two moles of aluminum we have 6 moles of HCl, 2 moles of AlCl3, and 3 moles of H2. If you take the mass you are given (78.4) and divide it by the molar mass (26.98), you will get the number of moles you have. The number I got out of my calculator was 2.90585618977. Can you do the rest or do you need more?

OpenStudy (lena772):

do i multiply that by hydrogens molar mass?

OpenStudy (anonymous):

Well you have 2.90585618977 moles of Aluminum, and for that you have 3 moles of hydrogen gas. \[\frac{ 2.90585618977 moles of Al}{ 1 } \times \frac{3 moles H2}{2 moles Al}\] then multiply across. \[\frac{ 8.71756856931}{ 2 }\] \[=4.35878428466 moles of H2\] Since the molar mass of H2 is 2, you'd multiply the amount of moles of H2 (4.35878428466) by 2. You end up getting 8.71756856931. In the original question you're limited by 3 significant figures, so you'd get 8.72 grams of H2.

OpenStudy (lena772):

So I guess I should go with b, since that's the closest

OpenStudy (kainui):

@Lena772 What you basically have is the relationship of how many atoms get shifted around into other atoms with that balanced equation. Literally all you do is rearrange to get from one side to the next like as if you're playing with legos or something. Hopefully I'm not wasting my time here, BUT whatever I think this should help in understanding the process of what's going on. Moles are just a way of counting things, just like a dozen eggs is just 12 eggs, a mol of atoms is just 602000000000000000000000 atoms. We also know that the periodic table tells us the amount of grams per mole of material we have. So let's combine this stuff together. We do this by just trying to divide out our unwanted units with the stuff we have until we get where we want to be. \[78.4 \ grams \ Al\] So what do we know of that divides this out? Well it's on the periodic table like I described above. \[78.4 \ grams \ Al * \frac{1 \ mol \ Al}{26.981 \ grams \ Al}\] See how now the grams of Al will divide out? We can treat units just like regular numbers! Just like 5/5=1 we have grams/grams=1. Now what's next? Well we have something that's in units of moles of aluminum. From the balanced equation you can see that you have 3 mols H2 per 2 mols of Al. Handy, that's the direction we're going for! \[78.4 \ grams \ Al * \frac{1 \ mol \ Al}{26.981 \ grams \ Al}*\frac{3 \ mol \ H_2}{2 \ mol \ Al}\] Notice I'm placing the fraction so that I always divide out the unit I want and then just looking for the number I need to go next to it when I need it. What's the next step to finish this problem off? We need to convert back to grams.

OpenStudy (lena772):

4.36 mol H2

OpenStudy (lena772):

4.36* 2= 8.72

OpenStudy (lena772):

but my only option is 8.79 so i should put that?

OpenStudy (lena772):

@Kainui

OpenStudy (kainui):

It's not unreasonable to expect your rounding error to be off by 7 hundredths of an amount when doing multiple conversions in a row. I'd say so.

OpenStudy (lena772):

thanks

OpenStudy (anonymous):

The error lies in the original molar mass of aluminum. Your teacher probably rounded it to 27.0 grams, not 26.98.

OpenStudy (lena772):

okay :)